OpenStudy (anonymous):

evaluate arcsec(2)? a) (pi/6, plus/minus 2pin, 5pi/6 plus/minus, 2pin b) (pie/3) c) (pie/6, 5pie/6) d) (pie/3 plus or minus 2pien, 5pie/3 plus or minus 2pien)

OpenStudy (blockcolder):

arcsec(2)=x sec(x)=2 cos(x)=1/2 x=arccos(1/2) Definitely A.

OpenStudy (lgbasallote):

or...input in a scientific calculator \(\mathsf{sec^{-1}x}\) i like that technique better

OpenStudy (lgbasallote):

just change that x to 2

OpenStudy (anonymous):

cos(x)=1/2 ---> x=pi/3

OpenStudy (anonymous):

@blockcorlder, your solution is correct but the choice is b.